# QUSL: Quantum Unsupervised Image Similarity Learning with Enhanced Performance

### v2

@@ Abstract
Leveraging quantum advantages to enhance machine learning capabilities has become a primary focus of research, particularly for complex tasks such as image similarity detection. To fully exploit the potential of quantum computing, it is essential to design quantum circuits tailored to the specific characteristics of the task at hand. In response to this challenge, we propose a novel quantum unsupervised similarity learning method — QUSL. Building upon the foundation of similarity detection triplets and generating positive samples through perturbations of anchor images, QUSL operates independently of classical oracles. By leveraging the performance of triplets and the characteristics of quantum circuits, QUSL systematically explores high-performance quantum circuit architectures customized for dataset features using metaheuristic algorithms, thereby achieving efficient quantum feature extraction with reduced circuit costs. Comprehensive numerical simulations and experiments on quantum computers demonstrate QUSL’s remarkable performance compared to state-of-the-art quantum methods. QUSL achieves reductions exceeding $50\%$ in critical quantum resource utilization while also realizing an enhancement of up to $19.5\%$ in similarity detection correlation across the DISC21, COCO, and landscape datasets. This enables efficient quantum similarity modeling for large-scale unlabeled image data with reduced quantum resource utilization.
@@ 1 Introduction
Quantum computing represents a paradigm shift in computational capabilities, harnessing the unique properties of quantum mechanics. Currently, quantum computers [1] have entered a stage known as Noisy Intermediate-Scale Quantum (NISQ) computing [2]. Don’t use a sledgehammer to crack a nut — quantum resources possess invaluable potential, and quantum computing demonstrates more apparent advantages when dealing with complex tasks [3, 4, 5, 6]. Quantum Machine Learning (QML) [7] has emerged as a noteworthy research focus in the field of image processing, giving rise to representative models such as quantum neural networks (QNN) [8, 9, 10], which have demonstrated remarkable performance in domains like quantum image classification [5, 11]. QML boasts several advantages [8, 12], one of which is its ability to utilize a larger feature space, providing enhanced capabilities for modeling complex patterns in image processing — a crucial aspect for handling high-dimensional tasks such as image processing.
@@
One particularly challenging task in the field of image processing is image similarity detection, which has far-reaching implications for complex applications such as target tracking and object recognition in real-world scenarios [13, 14]. It can be conceptualized as instances of image multiclassification problems characterized by a plethora of unknown classes, each containing only one member. Traditional approaches encounter prominent challenges. With increasing image sizes, the complexity and computational demands of image similarity detection escalate rapidly. This limits the feasibility of employing simple and intuitive algorithms[15, 16] (including Euclidean distance) in large-scale tasks, while the use of data dimensionality reduction techniques inevitably leads to information loss. This indirectly affects the generalization ability of unsupervised and self-supervised learning schemes [17, 18], resulting in unacceptable learning costs for large-scale, high-resolution image datasets.
To tackle these challenges, researchers have begun to explore the potential of quantum computing in image similarity tasks, leading to the development of the first wave of quantum similarity algorithms [19, 20] and quantum machine learning models [21, 22] specifically tailored to address challenges in image similarity analysis. Pioneering efforts like SliQ [23] have endeavored to establish unsupervised learning processes, where opportunities for enhancing both performance and learning strategies remain. However, a critical challenge in employing quantum computing for image similarity tasks lies in the design and training of high-performance quantum circuits that are tailored to the specific characteristics of the dataset.
@@
The design of quantum circuits presents a specialized optimization challenge [24, 25], aiming to minimize the usage of entanglement gates within the circuit structure’s shallowest possible depth while achieving near-equivalent functionality. Under such task scenarios, obtaining gradient information and ensuring differentiability are challenging, at times entirely unknown for quantum circuits involving intricate quantum state evolution processes. Consequently, addressing this issue through traditional gradient-based optimization methods proves to be notably challenging. Abstracting the optimization problem into a black-box scenario has emerged as an effective strategy to address this challenge [26]. As a result, the introduction of metaheuristic algorithms like evolutionary algorithms [27, 28, 29], which demonstrate self-organization and self-learning characteristics facilitating the design of more advanced quantum circuits [30], has become an efficient approach for quantum circuit design.
@@
In light of the aforementioned challenges and limitations, our research proposes a novel quantum unsupervised similarity learning framework — QUSL, which synergistically integrates the advantages of quantum computing and unsupervised learning. The core motivation behind QUSL is to fully harness the high expressive power of quantum circuits to more effectively process unsupervised image similarity tasks, breaking free from the reliance on classical image similarity algorithms and the performance constraints imposed by parameterized quantum circuit templates.
@@
In QUSL, we introduce an evolutionary algorithm-driven method to explore quantum circuit architectures customized for dataset features. This approach enables the automatic design of quantum circuits that are more efficient than fixed-template parameterized quantum circuits, effectively capturing the intricate patterns within image datasets and enhancing the performance of image similarity detection. Building upon this foundation, we propose a perturbation-based strategy for constructing similarity triplets, allowing QUSL to independently learn and detect image similarities without relying on classical algorithms, fully leveraging the advantages of quantum computing. QUSL provides a more generalizable and scalable solution for image similarity detection. The contributions of our research can be summarized as follows:
@@
    Numerical simulations and experiments on quantum computers across DISC21, COCO, and landscape datasets demonstrate QUSL’s superiority over state-of-the-art quantum methods, reducing critical quantum resource utilization by over $50\%$ while enhancing similarity detection correlation up by $19.5\%$, showcasing its performance advantage in near-realistic scenarios.
@@ 2.1 Qubits and Quantum Circuits
Applying a series of quantum gates to quantum bits yields a quantum circuit, and the design of any quantum computing task involves crafting a specific functional quantum circuit. NISQ quantum computers exhibit heightened sensitivity to larger circuit depths, thus emphasizing lower depths, minimal quantum gate counts, and more efficient utilization of quantum entanglement resources in quantum circuit design [31].
@@ 2.2 Quantum Machine Learning and Parametrized Quantum Circuit
QML is a method of performing machine learning tasks using quantum computers. In QML, parametrized quantum circuit (PQC) is a commonly used quantum circuit structure that can adapt to various types of scenes [32].
@@
Another prominent feature of PQC is that the learning process is reflected in the continuous adjustment of parameters, implying that quantum circuits need to be repeatedly executed with the same circuit depth a large number of times[33]. While some effective strategies can mitigate this issue [34], the accumulated total depth remains a key burden in large-scale learning tasks.
@@ 2.3 Heuristic Quantum Circuit Design
Evolutionary algorithms have emerged as powerful tools for automatically generating optimal quantum circuits tailored to specific datasets. The study of evolutionary strategies has matured significantly [27, 29, 35] finding widespread applications in diverse fields such as combinatorial optimization [36] and image quantum processing [37].
More specifically, evolutionary algorithms need to select a set of candidate quantum gates and appropriately describe the topology of the quantum circuit to adaptively design circuits. Both of these significantly influence the algorithm’s performance [38]. By strategically evolving, new individuals with higher fitness are obtained until the optimal quantum circuit is found or a predetermined endpoint is reached.
@@ 2.4 Unsupervised Quantum Approach for Image Similarity
Silver et al.’s groundbreaking work [23] introduced the first quantum learning approach for effective unsupervised image similarity detection on NISQ quantum computers. This method, rooted in classical unsupervised learning concepts such as Siamese and Triplet networks, utilizes variational quantum circuits for similarity learning and enhances model stability through a projection variance loss function, achieving promising performance on the landscape dataset.
However, the process of constructing triplets in SliQ relies on predefining the Euclidean distance to specify image similarity before training begins. From the perspective of leveraging quantum advantages, the introduction of classical image similarity severely restricts the potential advantages of this quantum model. On the other hand, SliQ employs a parametrized quantum circuit as the model ansatz and adopts a multi-layered hierarchical architecture to enhance its expressive power, yet it still fails to fully exploit the performance advantages of quantum circuits.
@@ 3.1 $A,P,N$-Triplet Construction and Quantum Embedding
$A,P,N$-triplets, a key concept in contrastive learning, have been effectively utilized for image similarity self-supervised learning [39, 40]. Each triplet comprises an anchor image, a positively labeled similar sample, and a negatively labeled dissimilar sample. During the learning process, these triplets are mapped to high-dimensional vectors in the embedding space, capturing the distance between vectors as a representation of similarity. However, unimproved $A,P,N$-triplets suffer from two drawbacks: inefficient mapping frequency and reliance on the oracle of visual concept discrimination [23]. The oracle, designed to preliminarily assess image similarity based on anchor images, selects positive and negative samples using algorithms or feature combinations. SliQ employs Euclidean distance as the oracle, incurring significant additional classical costs in quantum tasks involving large-scale, high-resolution unlabeled image datasets, thus constraining the quantum advantage.
QUSL use the method of feature interweaving generated by perturbing positive images to improve mapping efficiency and eliminate the use of the oracle of visual concept. Specifically, after anchor image selection is completed, similar images are obtained as positive images by applying slight perturbation to the anchor image as follow
@@
where, $I[c,i,j]$ denotes the anchor image with pixel index $(i,j)$ and RGB channel information $c$, $\overline{I}$ represents positive image, and $\mathcal{N}(0,\sigma^{2})$ signifies Gaussian noise with a mean of 00 and variance of $\sigma^{2}$. Constructing triplets in this stems from a simple fact that is easily overlooked: the most similar image to the anchor image is itself, and perturbation can bring slight differences to the anchor image to produce similarity differences. Gaussian noise, as a widely representative form of noise, exhibits stable interference effects on image features [41], making it a universal source of perturbation. In addition, QUSL adopts the strategy of utilizing feature interweaving to reduce the mapping process.
To adapt the quantum model to the task of RGB image similarity, classical images need to be quantum-embedded. Considering the limitations of the number of quantum bits in the NISQ era and the processing characteristics of quantum circuit models for image features, quantum amplitude embedding [42] is introduced to encode input features into physical quantum bits as follows
@@
where, $\left|\psi\right\rangle$ represents the quantum state obtained by amplitude embedding a pair of training set. For an image of size $N*N$, $p_{i}$ represents the $i$-th element of the training set vector, which is a normalized vector of length $6N^{2}$ after feature interweaving of the triplet. Thus, amplitude embedding saves a large number of quantum bits and preserves the complete image features mapped to the Hilbert space without the need for compression methods such as PCA. From the perspective of quantum information, feature interweaving of images in the triplet also results in complex entanglement between different image features, thereby augmenting the representation of image features [43]. The above process is shown in the Fig.1.
@@ 3.2 Dataset-Oriented Quantum Learning Circuit Construction
Unlike conventional triplets [39], the quantum version’s mapping procedure cannot be simply represented as the projection of three vectors onto a two-dimensional plane. This complexity stems from the incorporation of training sets, which introduces entanglement among image features, disturbing the direct correspondence between the mapped vectors and the features.
To reconcile this disparity, a corrective term must be included in the loss function for classical mapping to ensure projection coherence [23]. Based on the conceptualized quantum circuit model, the training sets of $A,P,N$-triplets are embedded into the Hilbert space and encapsulated by the probability distributions obtained from measuring the first four qubits of the quantum circuit across two mappings. By combining the practical significance of fitness and the rectified mapping relationship, the fitness function, designed to maintain projection consistency, is formulated as follows
@@
Performance symmetry arises from the principle of qubit permutation symmetry, where the numbering and arrangement of qubits do not affect the system’s evolution operator, maintaining system equivalence. This can lead to heterogeneous quantum circuits achieving similar or identical performance, hindering population diversity in evolutionary algorithms [44]. To mitigate this issue and maintain comprehensive coverage of the quantum circuit structure solution space, tournament selection can be employed, effectively reducing the survival rate of similar redundant individuals and balancing population diversity and convergence.
Performance constraints, another fundamental issue in quantum circuit design, require a holistic consideration of various aspects of quantum circuit performance evaluation. Non-dominated sorting [45] categorizes the population based on their support, with individuals within the same category being non-dominated. This approach generates diverse trade-off solutions, focusing on the Pareto front. When applied to quantum circuit design, non-dominated sorting simplifies population complexity and prevents the generation of high-depth, complex entangled quantum circuits. The synergistic application of tournament selection and non-dominated sorting ensures expansive exploration and expedites convergence to local depth. The comprehensive evolutionary strategy is illustrated in Fig. 2.
@@ 3.3 Image Similarity Performance Evaluation for Quantum Model
In the domain of unsupervised image similarity assessment and similar tasks, correlation analysis is frequently employed as a primary evaluation metric [46]. Correlation analysis evaluates model performance by assessing the consistency between the similarity scores generated by the model and the actual similarity between images, thereby addressing the limitations associated with directly comparing score disparities among different models. This classic comparative approach exhibits high versatility and can be directly transferred to the quantum domain.

### v4

@@ Abstract
Leveraging quantum properties to enhance complex learning tasks has been proven feasible, with excellent recent achievements in the field of unsupervised learning. However, current quantum schemes neglect adaptive adjustments for unsupervised task scenarios. This work proposes a novel quantum unsupervised similarity learning method — QUSL. Firstly, QUSL uses similarity triplets for unsupervised learning, generating positive samples by perturbing anchor images, achieving a learning process independent of classical algorithms. Subsequently, combining the feature interweaving of triplets, QUSL employs metaheuristic algorithms to systematically explore high- performance mapping processes, obtaining quantum circuit architectures more suitable for unsupervised image similarity tasks. Ultimately, QUSL realizes feature learning with lower quantum resource costs. Comprehensive numerical simulations and experiments on quantum computers demonstrate that QUSL outperforms state-of-the-art quantum methods. QUSL achieves over $50\%$ reduction in critical quantum resource utilization. QUSL improves similarity detection correlation by up to $19.5\%$ across multiple datasets, exhibiting robustness in NISQ environments. While using fewer quantum resources, QUSL shows potential for large-scale unsupervised tasks.
@@ 1 Introduction
Quantum computing represents a paradigm shift in computational capabilities and has entered a stage known as Noisy Intermediate-Scale Quantum (NISQ) computing Knill (2010); Preskill (2018). Quantum resources possess invaluable potential, and quantum computing demonstrates more apparent advantages when dealing with complex tasks Arute et al. (2019); Huang et al. (2022, 2021); Li et al. (2024). Quantum Machine Learning (QML) Biamonte et al. (2017) has emerged as a noteworthy research focus in the field of image processing, giving rise to representative models such as quantum neural networks (QNN) Schuld et al. (2014, 2015); Cong et al. (2019). QML boasts several advantages Schuld et al. (2014); Xu et al. (2021), one of which is its ability to utilize a larger feature space, providing enhanced capabilities for modeling complex patterns in image processing.
@@
One particularly challenging task in the field of image processing is image similarity detection, which has far-reaching implications for complex applications such as target tracking and object recognition in real-world scenarios Wang et al. (2023); Cheng et al. (2017). It can be conceptualized as instances of image multiclassification problems characterized by a plethora of unknown classes, each containing only one member. Classical approaches encounter prominent challenges. With increasing task scale, the complexity and computational demands of image similarity detection escalate rapidly. This limits the feasibility of employing simple and intuitive algorithms Zhang et al. (2011); Palubinskas (2017) in large-scale tasks, while the use of data dimensionality reduction techniques inevitably leads to information loss. This indirectly affects the generalization ability of unsupervised and self-supervised learning schemes Bai et al. ((2019); Wang et al. ((2014), resulting in unacceptable learning costs for large-scale, high-resolution image datasets.
Due to the potential of quantum computing in image processing tasks, the first wave of quantum similarity algorithms Dang et al. (2018); Liu et al. (2019) and quantum machine learning models Zhou and Sun (2015); Yan et al. ((2012) specifically tailored to address challenges in image similarity analysis have been developed. Series works have focused on establishing unsupervised learning processes by referencing classical models. Jaderberg et al. Jaderberg et al. (2022) adopted SimCLR’s unsupervised learning strategy Chen et al. (2020), enhancing the original network using QNNs. Silver et al.’s research Silver et al. ((2023) introduced the first quantum unsupervised image similarity model, employing tripletsMa et al. (2020) from contrastive learning to encode image pairs and likewise utilizing QNNs for feature extraction. Despite demonstrating promising performance, these works overlooked the necessity of adapting quantum learning circuits to suit specific task requirements. The critical challenge in employing quantum computing for image similarity tasks lies in the design and training of high-performance quantum circuits that are tailored to a specific task scenarios.
@@
The design of quantum circuits for learning scenarios presents a specialized optimization challenge Wang et al. (2022); Mohseni et al. (2022), aiming to minimize the usage of entanglement gates within the circuit structure’s shallowest possible depth while achieving superior expressive powerSim et al. (2019). Under such task scenarios, obtaining gradient information and ensuring differentiability are challenging, at times entirely unknown for quantum circuits involving intricate quantum state evolution processes. Consequently, the introduction of metaheuristic algorithms with self-organization and self-learning characteristics has become an efficient approach for obtaining more advanced quantum circuits Zhang and Zhao (2023); Ding and Spector ((2022); Krylov and Lukac ((2019); Rasconi and Oddi ((2019).
@@
In light of the aforementioned challenges and limitations, our research proposes a novel quantum unsupervised similarity learning framework — QUSL. The core motivation of QUSL is to fully harness the high expressive power of quantum circuits to more effectively process unsupervised image similarity tasks, breaking free from the reliance on classical image similarity algorithms and the performance constraints imposed by parameterized quantum circuit templates.
@@
QUSL introduce an evolutionary algorithm-driven method to explore quantum circuit architectures customized for dataset features. This approach enables the automatic design of quantum circuits that are more efficient than fixed-template parameterized quantum circuits, effectively capturing the intricate patterns within image datasets and enhancing the performance of image similarity detection. Subsequently, we propose a perturbation-based strategy for constructing similarity triplets, allowing QUSL to independently learn and detect image similarities without relying on classical algorithms, fully leveraging the advantages of quantum computing. QUSL provides a more generalizable and scalable solution for image similarity detection.
The highlights of our research can be summarized as follows:
@@
    Numerical simulations and experiments on quantum computers across five datasets demonstrate QUSL’s superiority over state-of-the-art quantum methods, reducing critical quantum resource utilization by over $50\%$ while enhancing similarity detection correlation up by $19.5\%$, showcasing its performance advantage in near-realistic scenarios.
@@ 2.1 Qubits and Quantum Circuits
Applying a series of quantum gates to quantum bits yields a quantum circuit, and the design of any quantum computing task involves crafting a specific functional quantum circuit. NISQ quantum computers exhibit heightened sensitivity to larger circuit depths, thus emphasizing lower depths, minimal quantum gate counts, and more efficient utilization of quantum entanglement resources in quantum circuit design Eisert (2021).
@@ 2.2 Quantum Machine Learning and Parametrized Quantum Circuit
QML is a method of performing machine learning tasks using quantum computers. In QML, parametrized quantum circuit (PQC) is a commonly used quantum circuit structure that can adapt to various types of scenes Benedetti et al. (2019).
@@
Another prominent feature of PQC is that the learning process is reflected in the continuous adjustment of parameters, implying that quantum circuits need to be repeatedly executed with the same circuit depth a large number of timesPérez-Salinas et al. (2020). While some effective strategies can mitigate this issue Cerezo et al. (2021), the accumulated total depth remains a key burden in large-scale learning tasks.
@@ 2.3 Heuristic Quantum Circuit Design
Evolutionary algorithms have emerged as powerful tools for automatically generating optimal quantum circuits tailored to specific datasets. The study of evolutionary strategies has matured significantly Zhang and Zhao (2023); Krylov and Lukac ((2019); Arufe et al. (2022) finding widespread applications in diverse fields such as combinatorial optimization Arufe et al. (2023) and image quantum processing Altares-López et al. (2021).
More specifically, evolutionary algorithms need to select a set of candidate quantum gates and appropriately describe the topology of the quantum circuit to adaptively design circuits. Both of these significantly influence the algorithm’s performance Wu et al. ((2023). By strategically evolving, new individuals with higher fitness are obtained until the optimal quantum circuit is found or a predetermined endpoint is reached.
@@ 2.4 Unsupervised Quantum Approach for Image Similarity
Silver et al.’s groundbreaking work Silver et al. ((2023) introduced the first quantum learning approach for effective unsupervised image similarity detection on NISQ quantum computers. SliQ utilizes triplets to encode image pairs, followed by similarity learning using variational quantum circuits, and enhances model stability through a projection variance loss function, achieving promising performance on the landscape dataset. SliQ adopted an early contrastive learning strategy, achieving a cost-effective transfer to the quantum domain.
However, the process of constructing triplets in SliQ relies on predefining the Euclidean distance to specify image similarity before training begins. On one hand, performing a large number of classical image similarity calculations severely limits the potential advantages of this quantum model. On the other hand, while SliQ employs a parametrized quantum circuit as the model ansatz and adopts a multi-layered hierarchical architecture to enhance its expressive power, it still fails to fully exploit the performance advantages of quantum circuits.
@@ 3.1 $A,P,N$-Triplet Construction and Quantum Embedding
$A,P,N$-triplets are a key concept in unsupervised learning, forming the basis for numerous mainstream self-supervised learning models Veit et al. ((2017); He et al. ((2020). Specifically, each triplet comprises an anchor image, a positively labeled similar sample, and a negatively labeled dissimilar sample. During the learning process, these triplets are mapped to high-dimensional vectors in the embedding space, learning sample features by capturing the distance between vectors.
SliQ employs triplets as a modeling process for classical image quantization, as illustrated in the lower part of Fig.1. In this process, SliQ uses Euclidean distance as an evaluation metric to select positive and negative samples. This approach incurs significant additional classical costs in quantum tasks involving large-scale, high-resolution unlabeled image datasets, thus limiting the method’s advantages.
QUSL uses the method of feature interweaving generated by perturbing positive images to improve mapping efficiency and eliminate the use of the oracle of visual concept. After anchor image selection, data augmentation is performed by applying slight perturbation to the anchor image as follow
@@
where $I[c,i,j]$ denotes the anchor image with pixel index $(i,j)$ and RGB channel information $c$, $\overline{I}$ represents the positive sample image, and $\mathcal{N}(0,\sigma^{2})$ signifies Gaussian noise with a mean of 00 and variance of $\sigma^{2}$. Gaussian noise, as a widely representative form of noise, exhibits stable interference effects on image features Hendrycks and Dietterich (2019), making it a universal source of perturbation. This method introduces similarity differences to the anchor image through noise perturbation, using these as positive samples. This approach draws inspiration from SimCLR’s data augmentation strategy Chen et al. (2020), greatly reducing the cost of constructing triplets. In addition, QUSL adopts the strategy of utilizing feature interweaving to reduce the mapping process.
In the encoding phase, QUSL uses quantum amplitude embedding Rebentrost et al. (2014) to encode input features into quantum bits as follows
@@
where, $\left|\psi\right\rangle$ represents the quantum state obtained by amplitude embedding a pair of training set. For an image of size $N*N$, $p_{i}$ represents the $i$-th element of the training set vector, which is a normalized vector of length $6N^{2}$ after feature interweaving of the triplet. Thus, amplitude embedding saves a large number of quantum bits and preserves the complete image features mapped to the Hilbert space without the need for compression methods such as PCA. From the perspective of quantum information, feature interweaving of images in the triplet also results in complex entanglement between different image features, thereby augmenting the representation of image features Paine et al. (2023). The above process is shown in the Fig.1.
@@ 3.2 Dataset-Oriented Quantum Learning Circuit Construction
Unlike conventional triplets Veit et al. ((2017), the quantum version’s mapping procedure cannot be simply represented as the projection of three vectors onto a two-dimensional plane. This complexity stems from the incorporation of training sets, which introduces entanglement among image features, disturbing the direct correspondence between the mapped vectors and the features.
To reconcile this disparity, a corrective term must be included in the loss function for classical mapping to ensure projection coherence Silver et al. ((2023). Based on the conceptualized quantum circuit model, the training sets of $A,P,N$-triplets are embedded into the Hilbert space and encapsulated by the probability distributions obtained from measuring the first four qubits of the quantum circuit across two mappings. By combining the practical significance of fitness and the rectified mapping relationship, the fitness function, designed to maintain projection consistency, is formulated as follows
@@
Performance symmetry arises from the principle of qubit permutation symmetry, where the numbering and arrangement of qubits do not affect the system’s evolution operator, maintaining system equivalence. This can lead to heterogeneous quantum circuits achieving similar or identical performance, hindering population diversity in evolutionary algorithms Nam et al. (2018). To mitigate this issue and maintain comprehensive coverage of the quantum circuit structure solution space, tournament selection can be employed, effectively reducing the survival rate of similar redundant individuals and balancing population diversity and convergence.
Performance constraints, another fundamental issue in quantum circuit design, require a holistic consideration of various aspects of quantum circuit performance evaluation. Non-dominated sorting Fang et al. (2008) categorizes the population based on their support, with individuals within the same category being non-dominated. This approach generates diverse trade-off solutions, focusing on the Pareto front. When applied to quantum circuit design, non-dominated sorting simplifies population complexity and prevents the generation of high-depth, complex entangled quantum circuits. The synergistic application of tournament selection and non-dominated sorting ensures expansive exploration and expedites convergence to local depth. The comprehensive evolutionary strategy is illustrated in Fig. 2.
@@ 3.3 Image Similarity Performance Evaluation for Quantum Model
In the domain of unsupervised image similarity assessment and similar tasks, correlation analysis is frequently employed as a primary evaluation metric Mei et al. (2023). Correlation analysis evaluates model performance by assessing the consistency between the similarity scores generated by the model and the actual similarity between images, thereby addressing the limitations associated with directly comparing score disparities among different models. This classic comparative approach exhibits high versatility and can be directly transferred to the quantum domain.
@@
The complete process of QUSL is shown in Fig. 3.
Figure 3: The complete process of QUSL. Some processes are summarized to ensure the clarity of the legend.